# Right for the Right Reasons: Avoiding Reasoning Shortcuts via Prototypical Neurosymbolic AI

### v1

@@ Abstract
Neurosymbolic AI is growing in popularity thanks to its ability to combine neural perception and symbolic reasoning in end-to-end trainable models. However, recent findings reveal these are prone to shortcut reasoning, i.e., to learning unindented concepts–or neural predicates–which exploit spurious correlations to satisfy the symbolic constraints. In this paper, we address reasoning shortcuts at their root cause and we introduce prototypical neurosymbolic architectures. These models are able to satisfy the symbolic constraints (be right) because they have learnt the correct basic concepts (for the right reasons) and not because of spurious correlations, even in extremely low data regimes. Leveraging the theory of prototypical learning, we demonstrate that we can effectively avoid reasoning shortcuts by training the models to satisfy the background knowledge while taking into account the similarity of the input with respect to the handful of labelled datapoints. We extensively validate our approach on the recently proposed rsbench benchmark suite in a variety of settings and tasks with very scarce supervision: we show significant improvements in learning the right concepts both in synthetic tasks (MNIST-EvenOdd and Kand-Logic) and real-world, high-stake ones (BDD-OIA). Our findings pave the way to prototype grounding as an effective, annotation-efficient strategy for safe and reliable neurosymbolic learning.  [^1] The code is available on our r4rr Github page.
@@ 2 Problem Statement
where $\mathcal{L}(p_{\theta},\mathcal{D},\textsc{K})=\frac{1}{|\mathcal{D}|}\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}}\log p_{\theta}(\mathbf{y}\mid\mathbf{x};K)$. Observe every RS corresponds to a deterministic optmimum for $\mathcal{L}$.
@@ 3 Prototypical Neurosymbolic AI
Given a finite dataset $\mathcal{D}$ and a concept $c\in\mathcal{C}=[h_{1}]\times\ldots\times[h_{k}]$, let
@@
be support set for concept $c$. This is the available set of datapoints which we can use to “anchor” the representation of each concept, hence avoiding the RS problem. For every concept $c\in\mathcal{C}$ we create a centroid or prototype $\mathbf{c}_{c}\in\mathbb{R}^{m}$ starting from $\mathcal{S}_{c}$ i.e., a representative vector where all datapoints belonging to concept $c$ should be mapped to. To this end, we use $k$ prototype extractors (one for each set of mutually exclusive concepts in $\mathcal{C}$). Each prototype extractor is an embedding function $f^{i}_{\mathbf{\theta}}:\mathbb{R}^{d}\to\mathbb{R}^{m_{i}}$ for $i\in[k]$ with learnable parameters $\theta$ which allows us to compute the centroids:
@@
We now relax the previous assumption that each concept must have at least one labelled datapoint, hence we can have $c\in\mathcal{C}$ such that $\mathcal{S}_{c}=\emptyset$. Instead, we require only that, within each set of mutually exclusive concepts (i.e., for every $h_{i}]$ with $i\in[k]$), there exist at least two datapoints labelled with two distinct concepts. This condition ensures that we can express the centroids of concepts without labelled examples as functions of the centroids of concepts with at least one labelled instance.
@@
where $\boldsymbol{\mu}_{\mathcal{H}_{i}}=\frac{1}{|\mathcal{H}|}\sum_{c^{\prime}\in\mathcal{H}}\mathbf{c}_{c^{\prime}}$ represents the mean of the known centroids in $h_{i}]$, and $\chi^{2}_{m,p}$ is the $p$-th lower-tail quantile of a $\chi^{2}$ distribution with $m$ degrees of freedom. Unless explicitly stated, the parameter $p$ is set to $0.99$ in our experiments. This ensures that, for each class $c$, the newly initialized centroid $\mathbf{c}_{c}$ has probability $0.99$ (over the random draw of $\epsilon$) of lying within the hyperball whose radius is given by the distance from $\boldsymbol{\mu}_{\mathcal{H}_{i}}$ to the farthest known centroid. This encourages separation among mutually exclusive classes, while reducing the risk that a zero-shot prototype is placed far from the region spanned by the known embeddings, which could otherwise lead to prototype collapse, where multiple concepts are mapped to a single, constraint-satisfying centroid.
@@
1:(i) Dataset $\mathcal{D}$, (ii) support set $\mathcal{S}_{c}$ for every concept $c\in\mathcal{C}$.
@@
13:   $\mathbf{c}_{c}\leftarrow\boldsymbol{\mu}_{\mathcal{H}}+\epsilon$ $\triangleright$ Compute centroids for unlabelled classes
14:  for $c$ in $\mathcal{E}$ do
15:for $(\mathbf{x},\mathbf{y})\in\mathcal{D}$ do
16:  $\mathcal{L}\leftarrow\mathcal{L}+\mathcal{L}^{\text{NeSy}}(\mathbf{y},\textsc{K})\triangleright$ If necessary for NeSy model, compute NeSy loss
@@ 4 Reasoning over Distances
where each $\nu:\bigcup_{i}{[h_{i}]}\to\{0,1\}$ represents (with a slight abuse of notation) a boolean assignment of the concepts in $\bigcup_{i}{[h_{i}]}$. We say that an assignment $\nu$ satisfies the knowledge K expressed as a propositional logic formula, denoted by $\nu\models\textsc{K}$, if and only if the recursive evaluation of formula K under assignment $\nu$ yields $1$. For every $c\in\bigcup_{i}{[h_{i}]}$ we define an independent Bernoulli random variable $Y_{c}\sim\text{Bernoulli}(y_{c})$. This induces a distribution over full truth assignments $\nu:\bigcup_{i}{[h_{i}]}\to\{0,1\}$. Then, we can rewrite the semantic loss as: $\mathcal{L}^{\text{NeSy}}(\mathbf{y},\textsc{K})=-\log\sum_{\nu\models\textsc{K}}p(\nu\mid\mathbf{y}).$
@@ 7 Related Work
NeSy encompasses a diverse set of approaches that integrate symbolic reasoning with neural networks [36, 17, 12, 14]. A central challenge in this field lies in reconciling discrete, non-differentiable symbolic reasoning with the gradient-based optimization underlying neural networks. To enable end-to-end training, many NeSy approaches soften the semantics of logical reasoning via probabilistic [26, 51, 23, 32, 1, 47, 41, 25, 48, 9] or fuzzy logic frameworks [10, 11, 45, 2, 40]. A common strategy involves compiling background knowledge into the loss function to encourage constraint satisfaction during training [49, 13, 2], though this typically offers no guarantee that constraints will hold at inference time. In contrast, methods that restructure the model’s output space can ensure constraint satisfaction by design [26, 15, 1, 7, 33, 18, 16, 19, 43, 44]. Despite the many advances [24, 46], [28, 31] showed that most NeSy systems are prone to RSs. Some mitigation strategies have been proposed [27, 31], but they can be either costly or they might fail in challenging settings like those proposed in rs-bench. To obviate to such problems, an interesting alternative direction has been taken by [30], where the authors do not try to obviate to the RSs problem but rather inform the user of when the model is taking a shortcut and facilitates acquiring informative dense annotations for mitigation purposes. For a comprehensive overview over the RSs problem in NeSy see [29].
@@ 60. Guidelines:
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
@@ Appendix D Tasks Details
The tasks belong to the rs-bench [^2] Link: https://github.com/unitn-sml/rsbench-code benchmark suite. In this benchmark, all ready-made data sets and generated datasets are distributed under the CC-BY-SA 4.0 license, with the exception of Kand-Logic, which is derived from Kandinsky-patterns [34] and as such is distributed under the GPL-3.0 license. Regarding the code, most of the code in the repo is distributed under the BSD 3 license, with the exception of Kand-Logic, which is derived from Kandinsky-patterns and as such is distributed under the GPL-3.0 license.
@@
Inspired by the artworks of Wassily Kandinsky, this task involves non-trivial perception to classify shapes and colors, and logical reasoning to infer patterns among them. Each datapoint $(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3})$ consists of three $64\times 64$ images, representing three geometric primitives each. An example is depicted in Figure 13. Every sample is labelled as 1 whether all images either include the same number of primitives with the same shape or with the same colour, and 0 otherwise. In order to formally specify the background knowledge K, we define six binary classes, deterministically determined from the concept classes:
@@
  • Single frames may lack contextual information, which can be important for understanding the scene. For example, in Figure 15(b) we can see that the context of the pedestrians is crucial for understanding the correct action to be taken. Indeed, even though in both frames the pedestrians are crossing the street, the context is different: in (a) the pedestrians are crossing the street and they are the only reason for the stop, while in (b) there is also a traffic light that is red. This means that in (a) the pedestrians are labeled as present because the car stops in front of them, while in (b) the pedestrians are not labeled as present because the car stops due to the red light.
@@ Appendix E Experimental Setup
To process the embeddings we created, the standard architecture of prototypical extractors from [42] was adapted to work with 1D convolutions. More in details, the prototypical architecture we employed adopts three convolutional blocks, including a $1\times 32$, $32\times 64$ and $64\times 128$ layer respectively, each comprising 5-wide convolutional filters, batch normalization, ReLU activation and a 1D kernel size $2$ max-pooling layer. The last block replaces max-pooling with adaptive average pooling of length 1. Finally, the output is flattened and further processed through a linear layer, 1D batch normalization and ReLU to yield a 256-dimensional embedding. For BDD-OIA, 21 prototypical extractors were employed, each classifying one concepts. As for the case of Kand-Logic, we modify the standard baseline to use the same number of backbones as the prototypical models.
@@
Likewise, for Kand-Logic the chosen transformations are scaling, anti-aliasing addition and translation, as the generated data are representative of the primitives the actual dataset contains. In the latter case, we generate 163 data points per (shape, colour) pair. For BDD-OIA, given a concept $j$, at least 6 samples are drawn with $i$ label 1. Then positive (resp. negative) samples are added from samples positive for each class $^{\prime}, jj$$\neq j^{\neq}$, where $j$ also occurs (resp. does not occur). We end up with $\approx 200$ samples per class, more than 85% of which are negatives. No data augmentation is used and the prototypical networks are all trained using one training batch randomly sampled at training time, with some classes being significantly under-represented, as shown in Table 7.
@@
All the experiments were run across ten seeds: 0, 128, 256, 512, 1024, 2048, 4096, 8192, 16384, 32768. The value of the following hyperparamters is used consistently for all the experiments when training the NeSy models to predict weak labels for each example in $\mathcal{D}$:
@@ Appendix F Additional Results
Since each concept is assigned a pair of different classes (i.e. a class of the shape and colour), distinguishing between inD and prD allows to effectively isolate the contribute of disentanglement from the one of prototypes. The results in Table 10 confirm the enhanced ability of PNet models in successfully retrieving the intended semantics for the intermediate concepts, while using them so comply with background knowledge. Moreover, they indicate that combining both forms of disentanglement with pretraining is generally the second best mitigation strategy across the board.
@@
In Table 8 we report an exhaustive picture for the comparison of shortcut mitigation strategies in MNIST-EvenOdd. The Table accounts for the mitigations introduced in [31] already discussed in the main text (Section 6 (RQ2)): these are partial concept supervision (C), input reconstruction (R) via the introduction of a reconstruction loss with an encoder-decoder architecture, Shannon entropy loss (H), and combinations thereof. The scores for baseline models are obtained from [5] and we used the code for all the mitigation strategies and combinations thereof as in [31]. Also, to convey the broader picture, we included the computation for concept (Acc(C)) and label (Acc(Y)) accuracy. As expected, we see that the results wrt Acc(C) mirror those reported in Table 4, where we reported the F1(C) for the same experiments. Moreover, our models consistently succeed in minimizing Cls(C), which is 0.0 across all the models considered. On the other hand, we see that the mitigation strategies generally do not have a great impact on the metrics over the final label (i.e., Acc(Y) and F1(Y)), as the NeSy models were already performing well on the unlabelled classes.
@@
Are prototypical NeSy models robust wrt the number of unlabelled datapoints?
In Figure 25(b) we report the results obtained in the same experiment reported in Section 6 (RQ3) wrt the metrics Acc(C), Acc(Y), F1(C) and F1(Y). Similarly to F1(C), we can see from the Figure (a) that the results we obtain for ACC(C) using our PNets are very stable and improve as we add more unsupervised data. On the contrary, as (b) shows, if we use the standard NeSy models taking advantage of the supervisions both in a pretraining stage (Pre) and at training time with a cross-entropy loss (C(S), the results present a high variance and are very susceptible to the percentage of unsupervised data used for training. Similar considerations hold for ACC(Y) as well in (c) and (d), with only the Semantic Loss baseline model maintaining label accuracy scores comparable with the PNet approach. Moreover, from (e) and (f) we notice the PNet approach consistently improves the standard NeSy models F1(Y) score as well (with SL being the only exception) for all the percentages of unsupervised data. Crucially, observe from (h) that all the standard models suffer from high concept collapse. How much each model collapses it predictions is again susceptible to the percentage of unsupervised data used for training. Now compare this behaviour with (g): our prototypical models have 0.0 concept collapse across all the percentages of unsupervised data.
@@
In Figure 25(c) and Figure 25(d) we report the results obtained in the experiments described in Section 6 (RQ4) wrt Acc(C), Acc(Y) and Cls(C) on the MNIST-EvenOdd task and the Kand-Logic task, respectively. As expected, the more classes are labelled the better the results become. However, it is interesting to note that—in the MNIST-EvenOdd task—even by labelling only two images we obtain better results over the concepts than using the standard NeSy models. The same happens for the Kand-Logic task, where we need at least two examples for the shape primitives and two for the colour primitives. However, we can again see that with just four labelled examples we get better results than all the standard NeSy models.
@@
In Figure 25(e) we report the results obtained wrt F1(C), Acc(C) and Cls(C) for different lower-tail quantiles of the $\chi^{2}$ distribution. Clearly, choosing $p=0.99$ improved our results—no matter the metric.

### v2

@@ Abstract
Neurosymbolic AI is growing in popularity thanks to its ability to combine neural perception and symbolic reasoning in end-to-end trainable models. However, recent findings reveal these are prone to shortcut reasoning, i.e., to learning unindented concepts–or neural predicates–which exploit spurious correlations to satisfy the symbolic constraints. In this paper, we address reasoning shortcuts at their root cause and we introduce Prototypical Neurosymbolic architectures. These models are able to satisfy the symbolic constraints (be right) because they have learnt the correct basic concepts (for the right reasons) and not because of spurious correlations, even in extremely low data regimes. Leveraging the theory of prototypical learning, we demonstrate that we can effectively avoid reasoning shortcuts by training the models to satisfy the background knowledge while taking into account the similarity of the input with respect to the handful of labelled datapoints. We extensively validate our approach on the recently proposed rsbench benchmark suite in a variety of settings and tasks with very scarce supervision: we show significant improvements in learning the right concepts both in synthetic tasks (MNIST-EvenOdd and Kand-Logic) and real-world, high-stake ones (BDD-OIA). Our findings pave the way to prototype grounding as an effective, annotation-efficient strategy for safe and reliable neurosymbolic learning.  [^1] The code is available on our r4rr Github page
@@ 2 Problem Statement
where $\mathcal{L}(p_{\theta},\mathcal{D},\textsc{K})=\frac{1}{|\mathcal{D}|}\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}}\log p_{\theta}(\mathbf{y}\mid\mathbf{x};K)$. Observe every RS corresponds to a deterministic optimum for $\mathcal{L}$.
@@ 3 Prototypical Neurosymbolic AI
Given a finite dataset $\mathcal{D}$ and a concept $c$, let
@@
be support set for concept $c$. This is the available set of datapoints which we can use to “anchor” the representation of each concept, hence avoiding the RS problem. For every concept $c$ we create a centroid or prototype $\mathbf{c}_{c}\in\mathbb{R}^{m}$ starting from $\mathcal{S}_{c}$ i.e., a representative vector where all datapoints belonging to concept $c$ should be mapped to. To this end, we use $k$ prototype extractors (one for each set of mutually exclusive concepts in $\mathcal{C}$). Each prototype extractor is an embedding function $f^{i}_{\mathbf{\theta}}:\mathbb{R}^{d}\to\mathbb{R}^{m_{i}}$ for $i\in[k]$ with learnable parameters $\theta$ which allows us to compute the centroids:
@@
We now relax the previous assumption that each concept must have at least one labelled datapoint, hence we can have a concept $c$ such that $\mathcal{S}_{c}=\emptyset$. Instead, we require only that, within each set of mutually exclusive concepts (i.e., for every $h_{i}]$ with $i\in[k]$), there exist at least two datapoints labelled with two distinct concepts. This condition ensures that we can express the centroids of concepts without labelled examples as functions of the centroids of concepts with at least one labelled instance.
@@
where $\boldsymbol{\mu}_{\mathcal{H}_{i}}=\frac{1}{|\mathcal{H}_{i}|}\sum_{c^{\prime}\in\mathcal{H}_{i}}\mathbf{c}_{c^{\prime}}$ represents the mean of the known centroids in $h_{i}]$, and $\chi^{2}_{m,p}$ is the $p$-th lower-tail quantile of a $\chi^{2}$ distribution with $m$ degrees of freedom. Unless explicitly stated, the parameter $p$ is set to $0.99$ in our experiments. This ensures that, for each class $c$, the newly initialized centroid $\mathbf{c}_{c}$ has probability $0.99$ (over the random draw of $\epsilon$) of lying within the hyperball whose radius is given by the distance from $\boldsymbol{\mu}_{\mathcal{H}_{i}}$ to the farthest known centroid. This encourages separation among mutually exclusive classes, while reducing the risk that a zero-shot prototype is placed far from the region spanned by the known embeddings, which could otherwise lead to prototype collapse, where multiple concepts are mapped to a single, constraint-satisfying centroid.
@@
1:(i) Dataset $\mathcal{D}$, (ii) support set $\mathcal{S}_{c}$ for every concept $c$.
@@
13:   $\mathbf{c}_{c}\leftarrow\boldsymbol{\mu}_{\mathcal{H}_{i}}+\epsilon$ $\triangleright$ Compute centroids for unlabelled classes
14:for $(\mathbf{x},\mathbf{y})\in\mathcal{D}$ do
15:  $\mathcal{L}\leftarrow\mathcal{L}+\mathcal{L}^{\text{NeSy}}(\mathbf{y},\textsc{K})\triangleright$ If necessary for NeSy model, compute NeSy loss
@@ 4 Reasoning over Distances
where each $\nu:\bigcup_{i}{[h_{i}]}\to\{0,1\}$ represents (with a slight abuse of notation) a boolean assignment of the concepts in $\bigcup_{i}{[h_{i}]}$. We say that an assignment $\nu$ satisfies the knowledge K expressed as a propositional logic formula, denoted by $\nu\models\textsc{K}$, if and only if the recursive evaluation of formula K under assignment $\nu$ yields 11. For every $c\in\bigcup_{i}{[h_{i}]}$ we define an independent Bernoulli random variable $Y_{c}\sim\text{Bernoulli}(y_{c})$. This induces a distribution over full truth assignments $\nu:\bigcup_{i}{[h_{i}]}\to\{0,1\}$. Then, we can rewrite the semantic loss as: $\mathcal{L}^{\text{NeSy}}(\mathbf{y},\textsc{K})=-\log\sum_{\nu\models\textsc{K}}p(\nu\mid\mathbf{y}).$
@@ 7 Related Work
NeSy encompasses a diverse set of approaches that integrate symbolic reasoning with neural networks [36, 17, 12, 14]. A central challenge in this field lies in reconciling discrete, non-differentiable symbolic reasoning with the gradient-based optimization underlying neural networks. To enable end-to-end training, many NeSy approaches soften the semantics of logical reasoning via probabilistic [26, 51, 23, 32, 1, 47, 41, 25, 48, 9] or fuzzy logic frameworks [10, 11, 45, 2, 40]. A common strategy involves compiling background knowledge into the loss function to encourage constraint satisfaction during training [49, 13, 2], though this typically offers no guarantee that constraints will hold at inference time. In contrast, methods that restructure the model’s output space can ensure constraint satisfaction by design [26, 15, 1, 7, 33, 18, 16, 19, 43, 44]. Despite the many advances [24, 46], [28, 31] showed that most NeSy systems are prone to RSs. Some mitigation strategies have been proposed [27, 31], but they can be either costly or they might fail in challenging settings like those proposed in rs-bench. To obviate to such problems, an interesting alternative direction has been taken by [30], where the authors do not try to avoid the RSs problem, but rather inform the user of when the model is taking a shortcut and facilitates acquiring informative dense annotations for mitigation purposes. For a comprehensive overview over the RSs problem in NeSy see [29].
@@ 60. Guidelines:
    • The name of the license (e.g., $C$-BY 4.0) should be included for each asset.
@@ Appendix D Tasks Details
The tasks belong to the rs-bench [^2] Link: https://github.com/unitn-sml/rsbench-code benchmark suite. In this benchmark, all ready-made data sets and generated datasets are distributed under the $C$-BY-SA 4.0 license, with the exception of Kand-Logic, which is derived from Kandinsky-patterns [34] and as such is distributed under the GPL-3.0 license. Regarding the code, most of the code in the repo is distributed under the BSD 3 license, with the exception of Kand-Logic, which is derived from Kandinsky-patterns and as such is distributed under the GPL-3.0 license.
@@
Inspired by the artworks of Wassily Kandinsky, this task involves non-trivial perception to classify shapes and colours, and logical reasoning to infer patterns among them. Each datapoint $(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3})$ consists of three $64\times 64$ images, representing three geometric primitives each. An example is depicted in Figure 13. Every sample is labelled as 1 whether all images either include the same number of primitives with the same shape or with the same colour, and 0 otherwise. In order to formally specify the background knowledge K, we define six binary classes, deterministically determined from the concept classes:
@@
  • Single frames may lack contextual information, which can be important for understanding the scene. For example, in Figure 15(b) we can see that the context of the pedestrians is crucial for understanding the correct action to be taken. Indeed, even though in both frames the pedestrians are crossing the street, the context is different: in (a) the pedestrians are crossing the street and they are the only reason for the stop, while in (b) there is also a traffic light that is red. This means that in (a) the pedestrians are labelled as present because the car stops in front of them, while in (b) the pedestrians are not labelled as present because the car stops due to the red light.
@@ Appendix E Experimental Setup
To process the embeddings we created, the standard architecture of prototypical extractors from [42] was adapted to work with 1D convolutions. More in details, the prototypical architecture we employed adopts three convolutional blocks, including a $1\times 32$, $32\times 64$ and $64\times 128$ layer respectively, each comprising 5-wide convolutional filters, batch normalization, ReLU activation and a 1D kernel size 22 max-pooling layer. The last block replaces max-pooling with adaptive average pooling of length 1. Finally, the output is flattened and further processed through a linear layer, 1D batch normalization and ReLU to yield a 256-dimensional embedding. For BDD-OIA, 21 prototypical extractors were employed, each classifying one concepts. As for the case of Kand-Logic, we modify the standard baseline to use the same number of backbones as the prototypical models.
@@
Likewise, for Kand-Logic the chosen transformations are scaling, anti-aliasing addition and translation, as the generated data are representative of the primitives the actual dataset contains. In the latter case, we generate 163 data points per (shape, colour) pair. For BDD-OIA, given a concept $j$, at least 6 samples are drawn with $i$ label 1. Then positive (resp. negative) samples are added from samples positive for each class $^{\prime}, jj$$\neq j^{\prime}$, where $j$ also occurs (resp. does not occur). We end up with $\approx 200$ samples per class, more than 85% of which are negatives. No data augmentation is used and the prototypical networks are all trained using one training batch randomly sampled at training time, with some classes being significantly under-represented, as shown in Table 7.
@@
All the experiments were run across ten seeds: 0, 128, 256, 512, 1024, 2048, 4096, 8192, 16384, 32768. The value of the following hyperparameters is used consistently for all the experiments when training the NeSy models to predict weak labels for each example in $\mathcal{D}$:
@@ Appendix F Additional Results
Since each concept is assigned a pair of different classes (i.e., a class of the shape and colour), distinguishing between inD and prD allows to effectively isolate the contribute of disentanglement from the one of prototypes. The results in Table 10 confirm the enhanced ability of PNet models in successfully retrieving the intended semantics for the intermediate concepts, while using them so comply with background knowledge. Moreover, they indicate that combining both forms of disentanglement with pretraining is generally the second best mitigation strategy across the board.
@@
In Table 8 we report an exhaustive picture for the comparison of shortcut mitigation strategies in MNIST-EvenOdd. The Table accounts for the mitigations introduced in [31] already discussed in the main text (Section 6 (RQ2)): these are partial concept supervision (C), input reconstruction (R) via the introduction of a reconstruction loss with an encoder-decoder architecture, Shannon entropy loss (H), and combinations thereof. The scores for baseline models are obtained from [5] and we used the code for all the mitigation strategies and combinations thereof as in [31]. Also, to convey the broader picture, we included the computation for concept (Acc(C)) and label (Acc(Y)) accuracy. As expected, we see that the results wrt. Acc(C) mirror those reported in Table 4, where we reported the F1(C) for the same experiments. Moreover, our models consistently succeed in minimizing Cls(C), which is 0.0 across all the models considered. On the other hand, we see that the mitigation strategies generally do not have a great impact on the metrics over the final label (i.e., Acc(Y) and F1(Y)), as the NeSy models were already performing well on the unlabelled classes.
@@
Are Prototypical NeSy models robust wrt. the number of unlabelled datapoints?
In Figure 25(b) we report the results obtained in the same experiment reported in Section 6 (RQ3) wrt. the metrics Acc(C), Acc(Y), F1(C) and F1(Y). Similarly to F1(C), we can see from the Figure (a) that the results we obtain for ACC(C) using our PNets are very stable and improve as we add more unsupervised data. On the contrary, as (b) shows, if we use the standard NeSy models taking advantage of the supervisions both in a pretraining stage (Pre) and at training time with a cross-entropy loss (C(S), the results present a high variance and are very susceptible to the percentage of unsupervised data used for training. Similar considerations hold for ACC(Y) as well in (c) and (d), with only the Semantic Loss baseline model maintaining label accuracy scores comparable with the PNet approach. Moreover, from (e) and (f) we notice the PNet approach consistently improves the standard NeSy models F1(Y) score as well (with SL being the only exception) for all the percentages of unsupervised data. Crucially, observe from (h) that all the standard models suffer from high concept collapse. How much each model collapses it predictions is again susceptible to the percentage of unsupervised data used for training. Now compare this behaviour with (g): our prototypical models have 0.0 concept collapse across all the percentages of unsupervised data.
@@
In Figure 25(c) and Figure 25(d) we report the results obtained in the experiments described in Section 6 (RQ4) wrt. Acc(C), Acc(Y) and Cls(C) on the MNIST-EvenOdd task and the Kand-Logic task, respectively. As expected, the more classes are labelled the better the results become. However, it is interesting to note that—in the MNIST-EvenOdd task—even by labelling only two images we obtain better results over the concepts than using the standard NeSy models. The same happens for the Kand-Logic task, where we need at least two examples for the shape primitives and two for the colour primitives. However, we can again see that with just four labelled examples we get better results than all the standard NeSy models.
@@
In Figure 25(e) we report the results obtained wrt. F1(C), Acc(C) and Cls(C) for different lower-tail quantiles of the $\chi^{2}$ distribution. Clearly, choosing $p=0.99$ improved our results—no matter the metric considered.